RC Meeting, Monday, October 18 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920

Accelerator Status

  1. Significant down time Friday swing into Saturday owl. After recovery (5:30 am Saturday), very efficient beam delivery the rest of the weekend.
  2. Pass change this morning (Hall A to 2 pass, Hall B to 1 pass)
  3. Maintenance day Wednesday
  4. Another pass change on Friday (4 pass)

Hall C Status

  1. So far we have collected about 3171 pions (as of 7:30 am) in this configuration
  2. Issues with bender LHe level Friday swing and Saturday owl - cryo adjustments seemed to mitigate for the rest of the weekend. Escorted Controlled Access (Heidi) to assess issue.

Current Run Plan

  1. Continue data taking until early Nov
  2. Escorted Controlled access Monday
  3. Longer access Wednesday

Current Issues

  1. Strange issue with HV controls (SHMS 2Y11- powers off when bringing up the group) [1]. Doesn't seem to happen with new HV gui: Update: this issues seems to have gone away
  2. Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
  3. The HB has seen some fluctuations in the LHe levels. At the next opportunistic access Heidi would like to investigate the cryo line. ( PENDING )
    • Heidi believes this is related to a vacuum issue and the access is for diagnostic purposes only. Can be escorted during regular office hours (do not call her in at midnight for this)
    • If indeed it proves to be the case the only remedy would be to warm up the magnet and pump the region (which would be a multiple day operation)
  4. Water flow to the dump aperture needs to be replaced: needs 4+h and Steve Suhring believes that it can wait until the SAD ( PENDING )
  5. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.
  6. Still running with no UPS on the SHMS cryo controls - Steve L. will put in PR for batteries and UPS. Wants to inspect existing system at next opportunity. ( PENDING )
  7. HCCAM02 not functioning (lift required to replace).

Special Requests

Useful Links